# Large Language Models (LLMs) for Wireless Networks: An Overview from the Prompt Engineering Perspective

### v1

@@ Abstract
Large language models (LLMs) have been successfully applied to many fields, showing outstanding comprehension and reasoning capabilities. Despite their great potential, LLMs usually require dedicated fine-tuning for domain-specific applications such as wireless networks. Such domain adaptation can be extremely demanding for computational resources, while many network devices have limited computation capacity. To this end, this work explores LLM-enabled wireless networks from the prompt engineering perspective, i.e., designing input prompts to guide LLMs to generate desired output. Compared with other approaches to using LLMs, prompt engineering is more flexible and resource-efficient, which better aligns with wireless network features. In particular, we first introduce LLM fundamentals and compare different prompting techniques such as in-context learning, chain-of-thought, and self-refinement. Then we propose two novel prompting schemes for network applications, namely iterative prompting for wireless network optimization, and self-refined prompting for network prediction. The case study shows that the proposed schemes can achieve comparable performance as conventional machine learning techniques, and our prompting-based methods avoid the complexity of dedicated model training and fine-tuning, which is one of the key bottlenecks of existing machine learning techniques.
@@ I Introduction
Approaches & Steps Main features Advantages Potential difficulties Computational resources & Time costs Possible wireless network applications
@@
Pre-training LLMs Training LLMs from scratch on large-scale datasets, typically involving hundreds of billions of tokens for the next token prediction. It is a foundational step of LLM development and usage, including dataset collection and model training. Pre-training enables LLMs fundamental and crucial capabilities, such as comprehension, reasoning, zero-shot generation, and instruction following, which can address various downstream tasks. Pre-training is extremely resource-intensive, requiring massive labelled data, long training times, substantial energy consumption, and a tremendous amount of hardware accelerators like GPUs or TPUs. High Thermal Design Power GPUs such NVIDIA H100100100100-$80$G or A100100100100-$80$G are required. For example, training Llama3.13.13.13.1-$405$B took approximately $30.84$M GPU hours with $15$T tokens from publicly available sources and $25$M synthetic data. A network-specific LLM indicates great benefits, but training a network-specific LLM from scratch may be inappropriate due to the computational resource requirements.
@@
Fine-tuning Fine-tuning refers to adapting a pre-trained LLM to specific tasks by partially or completely updating the model weights or introducing additional adapter modules to fit smaller and domain-specific datasets. Fine-tuning significantly reduces time and computational cost compared to pre-training, producing LLM models tailored to specific tasks. Fine-tuning has to consider the risk of overfitting and losing parts of the generalizability. Labelled data or pairwise comparison data may be needed, involving additional human labour force. The cost may vary between model sizes and fine-tuning methods. Typically, if full fine-tuning a 7B parameter model with 8 A100100100100-$80$G GPUs, it could be $100$-$200$ GPU hours on a $100$M token dataset (assuming a maximum length of $512$ tokens). Fine-tuning is a realistic approach to apply LLMs to wireless networks, adapting a general-domain LLM to specific tasks such as network troubleshooting, project coding, and network configuration.
@@
Retrieval augmented generation (RAG) RAG combines LLMs with an external knowledge base, enabling the model to retrieve relevant information during inference. It enhances LLM’s capability to provide accurate and up-to-date responses, particularly in knowledge-intensive tasks. RAG can improve LLM’s response accuracy by leveraging external knowledge sources. It is particularly useful when the model’s training data is outdated or lacks specific information. RAG involves extra integration complexity, as it requires efficient retrieval algorithms since the inference latency may increase if the retrieval process is not optimized. Moreover, RAG highly depends on a well-structured knowledge base, which cannot be easily created and maintained. The computational cost depends on the retrieval method and the size of the knowledge base. Inference may take seconds to minutes per query, depending on the size of the base language model. Requires a balance between retrieval speed and accuracy, which can influence the choice of hardware, e.g., using faster storage for the knowledge base. Considering the cost of fine-tuning, RAG is a more practical approach for implementing LLMs, especially when existing network-specific datasets are available. It ensures the LLM has access to the most relevant and up-to-date network data, improving its accuracy for addressing the latest network tasks.
@@
Prompt engineering Prompting engineering refers to crafting specific input prompts to guide a pre-trained LLM to generate desired outputs. It leverages the existing capabilities of the LLM without any model retraining or tuning. Prompt engineering has low computational costs and quick implementation. It is also flexible and can be easily adapted to different tasks without additional training. It requires expertise in crafting effective prompts, as the quality of the output heavily depends on how the prompt is formulated. It is also limited by the inherent capabilities of the pre-trained LLMs. It has minimal computational costs and often requires just a single GPU or a few GPUs. Crafting and experimenting with different prompts may take only a few hours. Time costs are significantly lower compared to other methods, making it suitable for rapid prototyping and testing. Prompt engineering enables rapid decision-making and responses in dynamic wireless network environments. It aligns well with many network tasks, especially considering the limited computational resources of network devices.
@@
    Note that these approaches can be combined to achieve better performance, e.g., fine-tuning network LLMs and then prompting.
@@
As a sub-field of generative AI, large language models (LLMs) have received considerable interest from industry and academia [1]. After being trained on a large number of datasets, LLMs have shown revolutionary comprehension and reasoning capabilities. The advancement of generative AI and LLMs also provides promising opportunities for 6G wireless networks, including integrated satellite-aerial-terrestrial networks [2], vehicular networks [3], semantic communication [4], etc. Despite the great potential, integrating LLMs into wireless networks still faces several challenges. Firstly, wireless networks are complex large-scale systems with various knowledge domains, i.e., signal processing and transmission, network architecture and design, protocol, standards, and so on. Applying general-domain LLMs directly to domain-specific network tasks may lead to poor performance. Secondly, LLMs are extremely demanding for computational resources. LLM pre-training and fine-tuning are usually implemented on high-performance GPUs such as NVIDIA A100 and H100. However, wireless network devices usually have limited computational and storage capacities, which may prevent practical LLM network applications. In addition, LLMs can be applied in various approaches, including pre-training models from scratch, fine-tuning LLMs for domain-specific tasks, retrieval augmented generation (RAG), prompt engineering, etc. Each method has unique features and potential difficulties, and it is critical to identify the most efficient methods to adapt LLMs to wireless networks.
@@
Given the above opportunities and challenges, this work focuses on prompt engineering, which is regarded as one of the most resource-efficient approaches to using LLMs. In particular, prompting refers to designing input prompts to guide pre-trained LLMs to generate desired outputs. It takes advantage of the inherent inference capabilities of pre-trained LLMs, and avoids the complexity of dedicated model training or fine-tuning. Compared with other approaches, prompting-based techniques have several key advantages: 1) Resource-efficient: Compared with pre-training and fine-tuning, prompt engineering techniques have much lower requirements on computational resources like a single GPU. Such a resource-efficient approach can be easily deployed in many network devices. 2) Higher flexibility: Prompting-based methods can quickly adapt to various tasks by crafting the corresponding demonstrations and queries. It indicates an efficient method to customize LLMs to address a wide range of network tasks. 3) Fast implementation: Prompt engineering relies on the inference capabilities of LLMs, and it avoids the time cost of updating LLM parameters. Therefore, the low response time can better handle low-latency services in wireless networks.
@@
LLMs have been discussed in several existing studies, but they mainly focus on system-level discussions and module designs, e.g., edge intelligence [5], grounding and alignment[6], 6G security[7], and our previous work also investigates LLM-enabled power control and traffic prediction [8, 9]. However, this work is different from existing studies by systematically exploring prompt engineering and wireless network applications, providing detailed prompt designs and specific case studies. The main contributions include:
@@
1) Firstly, it presents in-depth analyses of LLM fundamentals such as pre-training, fine-tuning, RAG, and prompt engineering. We compare these techniques regarding main features, computational resources and time costs, potential difficulties, and wireless network applications. Then, it overviews multiple prompting techniques, such as in-context learning, chain-of-thought, prompt-based planning, and self-refinement, and we further investigate how various prompting techniques can adapt to wireless networks in terms of main features, advantages, and potential issues.
@@
2) We propose two novel prompting techniques, namely iterative prompting and self-refined prompting, aiming to address network optimization and prediction problems, respectively. Specifically, iterative prompting can learn from previous experience and exploration to improve LLM’s performance on target tasks iteratively, which is mainly designed for network optimization problems. By contrast, self-refined prompting allows them to correct their outputs through iterative feedback and refinement demonstration prompts, aiming to address network prediction tasks. These schemes rely on LLM’s inference capabilities, and such resource-efficient techniques align well with the limited computational capacities of network devices. In addition, our proposed algorithms can be easily generalized to various wireless network applications, and the case studies show that they achieve satisfactory performance in network power control and traffic prediction problems.
@@ II LLM Fundamentals towards Wireless Network Applications
This section introduces various approaches for applying LLMs to wireless networks, including pre-training, fine-tuning, RAG, and prompt engineering. Specifically, Table I compares these techniques in terms of main features, advantages, potential difficulties, computational resources and time costs, and possible wireless network applications.
@@
Firstly, pre-training is a foundational step for LLM development, which includes data collection, filtering, and model training. Pre-training enables LLMs to develop fundamental capabilities such as comprehension, reasoning, and instruction following. However, it requires considerable computational resources such as NVIDIA H100100100100-$80$G or A100100100100-$80$G for implementation, and the training process may take weeks and months. Therefore, despite the great benefits, pre-training a wireless-specific LLM from scratch can be extremely time-consuming with huge costs of computational resources.
@@
By contrast, fine-tuning is a more affordable approach for LLM implementation. In particular, fine-tuning indicates adapting a pre-trained general-domain LLM to specific tasks by partially updating the model weights to fit smaller and domain-specific datasets such as 3GPP standards. The fine-tuning costs are related to the LLM model size and fine-tuning methods. For instance, fully fine-tuning a 7B LLM model using 8 A100 GPUs may take 100-200 GPU hours for a 100M token dataset, e.g., SPEC5G dataset with 134M words[10] and 3GPP dataset Tspec-LLM with 534M words[11]. Similarly, RAG is another promising direction for practical LLM applications for communication networks. It integrates LLMs with an external knowledge base, and then the LLM can retrieve related information from the external knowledge source during the inference. RAG can significantly improve the quality and accuracy of the generated content, since the LLM has access to the most relevant and up-to-date information.
Compared with previous techniques, prompt engineering has much lower computational costs and faster implementation. Instead of updating model parameters, it relies on LLM’s inherent capabilities, designing specific input prompts to guide the generation of pre-trained LLMs. Without the model updating burdens, prompt design and testing only need a few GPUs, and the time costs are significantly lower than fine-tuning or RAG. Due to its low costs, it can also quickly adapt to dynamic network environments, handling various network tasks with much lower response time.
@@
The above analyses show that prompt engineering is one of the most efficient approaches to using LLMs, which aligns well with several crucial wireless network features. Firstly, many network devices are computational and energy resource-constrained, and prompt engineering has low requirements for computational resources, which can be easily deployed and implemented. Secondly, prompt engineering allows human language-based instructions, e.g., network-specific task descriptions, questions, and solutions. By integrating natural language, network intelligence will become more accessible, especially for operators without professional AI knowledge. In addition, it has higher design flexibility. With proper demonstrations, prompts can be quickly designed and applied to various network tasks, making rapid responses to network dynamics. Therefore, it is crucial to investigate the application of prompting techniques to wireless networks, showing a promising way to utilize generative AI.
@@
TABLE II: Summary of different prompting techniques for wireless networks.
@@
Prompting techniques Main features Advantages Potential issues Wireless network applications
@@
In-context learning In-context learning empowers LLMs to learn from contextual information provided in task-specific instruction and demonstrations. These prompts are designed with relevant, task-specific examples that are converted into natural language comprehensible to LLM. LLMs can be generalized to perform various learning tasks by utilizing existing knowledge from pre-training data and acquiring new task-solving strategies from contextual demonstration prompts. In-context learning is a fundamental and crucial capability of LLMs. Demonstrations greatly impact in-context learning, which should be carefully designed, including the selection, format, and order of examples, and then transform the selected examples into a formatted prompt. In-context learning is a promising approach to applying LLMs to wireless networks, using previously accumulated solutions as input prompts to address new network issues.
@@
Chain-of-thought Chain-of-thought is an advanced strategy to improve LLM’s capabilities in complex reasoning, which incorporates sequences of intermediate reasoning steps into demonstration prompts. It can guide LLMs to better understand the logical steps from questions to answers. LLMs can solve complex problems more transparently and logically through chain-of-thought demonstration prompts, such as “Let’s think step by step”. Existing studies have shown that chain-of-thought can significantly improve LLM’s performance in diverse tasks. Despite the advancement, chain-of-thought prompting still faces issues like incorrect reasoning and instability. Although LLMs generate correct final answers, their inherent instability may provide incorrect and misleading intermediate reasoning steps. The step-by-step chain-of-thought prompts can make LLMs better adapt to complicated wireless environments, which may require step-by-step reasoning, planning, and decision-making, e.g., project code generation with multi-step scheduling.
@@
Prompt-based planning Different from in-context learning and chain-of-thought, prompt-based planning aims to address more complicated scenarios such as multi-hop question answering. It breaks complex tasks down into manageable sub-tasks. In-context learning and chain-of-thought are more suitable for single-task implementation. By contrast, LLMs can make complex tasks more manageable by using their planning capabilities, lowering the overall difficulties. Currently, LLM-based planning requires dedicated analyses to decompose a complicated task into multiple sub-tasks. Zero-shoot and automated task decomposition is still the main difficulty. Wireless networks are complicated systems and planning is a crucial capability for addressing many network tasks. LLM-enabled automated task decomposition can be extremely useful for long-term network management and project development.
@@
Self-refinement Considering that LLMs could initially generate incorrect answers, self-refinement allows them to correct their outputs through iterative feedback and refinement demonstration prompts. Through iterative feedback generation and prediction refinement, LLMs can self-enhance their outputs with strong effectiveness and scalable capabilities. To avoid repeating previous mistakes, all historical outputs are iteratively appended as the inputs of LLMs. It may lead to increasingly lengthy demonstration prompts and unavoidable resource consumption. Self-refinement will significantly save human efforts when using LLMs to address network tasks, since LLMs can constantly improve the previous network policy automatically.

### v2

# Large Language Models for Wireless Networks: An Overview from the Prompt Engineering Perspective

@@ Abstract
Recently, large language models (LLMs) have been successfully applied to many fields, showing outstanding comprehension and reasoning capabilities. Despite their great potential, LLMs usually require dedicated pre-training and fine-tuning for domain-specific applications such as wireless networks. These adaptations can be extremely demanding for computational resources and datasets, while most network devices have limited computation power, and there are a limited number of high-quality networking datasets. To this end, this work explores LLM-enabled wireless networks from the prompt engineering perspective, i.e., designing prompts to guide LLMs to generate desired output without updating LLM parameters. Compared with other LLM-driven methods, prompt engineering can better align with the demands of wireless network devices, e.g., higher deployment flexibility, rapid response time, and lower requirements on computation power. In particular, this work first introduces LLM fundamentals and compares different prompting techniques such as in-context learning, chain-of-thought, and self-refinement. Then we propose two novel prompting schemes for network applications: iterative prompting for network optimization, and self-refined prompting for network prediction. The case studies show that the proposed schemes can achieve comparable performance as conventional machine learning techniques, and our proposed prompting-based methods avoid the complexity of dedicated model training and fine-tuning, which is one of the key bottlenecks of existing machine learning techniques.
@@ I Introduction
Approaches Key principles Advantages Potential difficulties Computational resources & Time costs Possible wireless network applications
@@
Pre-training LLMs 1) Training LLMs from scratch on hundreds of billions of tokens for the next token prediction.   2) Pre-training is a foundational step of LLM development and usage. Pre-training enables LLMs with comprehension, reasoning, and instruction following capabilities, which can address various downstream tasks. Pre-training is extremely resource-intensive, requiring long training times, and substantial energy consumption and GPUs or TPUs. High Thermal Design Power GPUs such NVIDIA H100100100100-$80$G or A100100100100-$80$G are required, e.g., training Llama3.13.13.13.1-$405$B took approximately $30.84$M GPU hours. A network-specific LLM indicates great benefits, but training network LLMs from scratch may be inappropriate due to the computational resource requirements.
@@
Fine-tuning Fine-tuning refers to adapting a pre-trained LLM to specific tasks by partially or completely updating the model weights or introducing additional adapter modules. Fine-tuning significantly reduces time and computational cost compared to pre-training, producing LLM models tailored to specific tasks. 1) Fine-tuning has to consider the risk of overfitting and losing generalization.     2) Labelled data may be needed with additional human labour. The cost may vary between model sizes and fine-tuning methods. If full fine-tuning a 7B parameter model with 8 A100100100100-$80$G GPUs, it could be $100$-$200$ GPU hours on a $100$M token dataset. Fine-tuning is a realistic approach to apply LLMs to wireless networks, adapting a general-domain LLM to specific tasks, e.g., network troubleshooting and configuration.
@@
Retrieval augmented generation (RAG) RAG combines LLMs with an external knowledge base, enabling the model to retrieve relevant information during inference and provide more accurate responses. RAG can improve LLM’s response accuracy by leveraging external knowledge sources, generating more up-to-date results. 1) RAG involves extra integration complexity, as it requires efficient retrieval algorithms.   2) RAG highly depends on a well-structured knowledge base, requiring dedicated creation. 1) The cost depends on the retrieval method and the size of the knowledge base. Inference may take seconds to minutes per query.     2) Requires a balance between retrieval speed and accuracy. 1) RAG is a more practical approach for implementing LLMs, especially when existing network-specific datasets are available.    2) It ensures the LLM has access to the most relevant and up-to-date network data.
@@
Prompt engineering Prompt engineering refers to crafting specific inputs to guide a pre-trained LLM to generate desired outputs, without any model retraining or tuning. 1) Prompt engineering has low computational costs and quick implementation. 2) It can be easily adapted to different tasks without additional training. 1) It requires expertise in crafting effective prompts. 2) It is also limited by the inherent capabilities and context window size of the pre-trained LLMs. 1) Without the need for backward passes and gradient updates, it requires minimal computational resources.   2) Time costs are significantly lower compared to other methods. 1) Prompt engineering enables rapid decision-making and responses in dynamic wireless network environments.    2) It aligns well with many network tasks.
@@
    Note that the objective of this table is to introduce the features of each technique, and they serve different purposes in LLM development and application.
@@
    These approaches can be combined to achieve better performance, e.g., fine-tuning network LLMs and then prompting.
@@
As a sub-field of generative AI, large language models (LLMs) have received considerable interest from industry and academia [1]. The advancement of generative AI and LLMs also provides promising opportunities for 6G networks, including strong reasoning and planning capabilities, multi-modal understanding for 6G sensing, semantic communication [2], integrated satellite-aerial-terrestrial networks [3], etc. Despite the great potential, integrating LLMs into wireless networks still faces several challenges. Firstly, wireless networks are complex large-scale systems with various knowledge domains, i.e., signal processing and transmission, network architecture and design, protocol, standards, and so on. Applying general-domain LLMs directly to domain-specific network tasks may lead to poor performance. Secondly, the development of LLMs relies on high-quality datasets for fine-tuning adaptation, while there is a limited amount of high-quality networking datasets such as SPEC5G and Tspec-LLM [1]. Moreover, LLMs are extremely demanding in terms of computational resources. LLM pre-training and fine-tuning are usually implemented on high-performance GPUs such as NVIDIA A100 and H100, but wireless network devices usually have limited computational and storage capacities. LLMs involve a broad range of techniques, such as pre-training, fine-tuning LLMs for domain-specific tasks, retrieval augmented generation (RAG), prompt engineering, etc. Therefore, it is critical to identify an efficient method to better adapt LLMs to wireless networks.
@@
Given the above opportunities and challenges, this work introduces prompt engineering, which is regarded as a resource-efficient and flexible approach to using LLMs with fast implementation speed[4]. These advantages will help to overcome the above LLM application challenges, e.g., deployment difficulties and requirements for computational resources. In particular, prompting refers to designing input prompts to guide pre-trained LLMs to generate desired outputs. It takes advantage of the inherent inference capabilities of pre-trained LLMs, and avoids the need for backward passes and gradient updates. Therefore, prompt engineering has several key features: 1) Resource-efficient: Prompting only needs forward passing of the model, and no need to store all intermediate activations for backpropagation. Such a resource-efficient approach may mitigate the computational burden of network servers and devices. 2) Higher flexibility: Prompting-based methods can quickly adapt to various tasks by crafting the corresponding textual demonstrations and queries without extra coding steps[5]. It indicates an efficient method to customize LLMs to address a wide range of network tasks. 3) Fast implementation: Prompt engineering relies on the inference capabilities of LLMs, and it avoids the time cost of updating LLM parameters. Therefore, the low response time can better handle low-latency network services. Finally, note that although prompting LLMs can significantly enhance the performance, it is not a standalone solution for LLM usage[6]. For instance, combining prompt engineering with fine-tuning may create a robust framework with both prompting flexibility and model specialization. However, this work focuses on prompting techniques to investigate their full potential.
@@
LLMs have been discussed in several existing studies, but they mainly focus on system-level discussions and module designs, e.g., edge intelligence [7] and grounding and alignment[8]. Prompting engineering is also used for AI-generated everything services in [9], and our previous work also investigates LLM-enabled power control and traffic prediction in [10, 11]. However, this work is different from existing studies by systematically exploring prompt engineering and wireless network applications, providing detailed prompt designs and specific case studies. The main contributions are:
@@
1) Firstly, we present in-depth analyses of LLM fundamentals and the feasibility of network applications such as pre-training, fine-tuning, RAG, and prompt engineering. Each technique is introduced in terms of computational resources and time costs, potential difficulties, and wireless network applications. Then, we provide an overview of a variety of prompting techniques and discuss their applications to wireless networks, such as in-context learning, chain-of-thought, prompt-based planning, and self-refinement.
@@
2) We propose two novel prompting techniques, namely iterative prompting and self-refined prompting, aiming to address network optimization and prediction problems, respectively. Specifically, iterative prompting learns from previous experiences to improve LLM’s performance on target tasks iteratively, which is suitable for handling network optimization problems. By contrast, self-refined prompting can correct their outputs through iterative feedback and refinement prompts, aiming to address network prediction tasks. These schemes rely on LLM’s inference capabilities, and such resource-efficient techniques align well with the limited computational capacities of many network devices. In addition, our proposed algorithms can be easily generalized to various wireless network applications, and the case studies show that they achieve satisfactory performance in network power control and traffic prediction problems.
Figure 1: Illustration of different prompting techniques.
@@ II LLM Fundamentals towards Wireless Network Applications
This section introduces various approaches for integrating LLMs to wireless networks, including pre-training, fine-tuning, RAG, and prompt engineering. Specifically, Table I presents these techniques in terms of their working principle, advantages, potential difficulties, computational resources and time costs, and possible wireless network applications.
@@
Firstly, pre-training is a foundational step for LLM development, which includes data collection, filtering, and model training. Pre-training enables LLMs to develop fundamental capabilities such as comprehension, reasoning, and instruction following. Then, these capabilities are the foundation for developing other LLM-inspired techniques such as prompt engineering and planning, e.g., prompt engineering relies on the inherent capabilities of LLMs. However, it requires considerable computational resources such as NVIDIA H100100100100-$80$G or A100100100100-$80$G for implementation, and the training process may take weeks and months.
@@
Given the training costs, pre-training a wireless-specific LLM from scratch can be extremely time and energy-consuming. By contrast, fine-tuning is a more affordable approach for adapting LLMs to network domains. Given pre-trained general-domain LLMs, fine-tuning indicates adapting LLMs to specific tasks by partially updating the model weights to fit smaller and domain-specific datasets such as 3GPP, 5G, and O-RAN standards. The fine-tuning costs are related to the LLM model size and fine-tuning methods. For instance, fully fine-tuning a 7B LLM model using 8 A100 GPUs may take 100-200 GPU hours for a 100M token dataset. For large-scale datasets in wireless fields, e.g., SPEC5G dataset with 134M words and 3GPP dataset Tspec-LLM with 534M words [10], the costs on computational resources may be larger. Similarly, RAG is another promising direction for practical LLM applications for wireless networks. It integrates LLMs with an external knowledge base, and then the LLM can retrieve related information from the external knowledge source during the inference. Since LLMs have access to the most relevant and up-to-date information, e.g., novel network standards, architecture and signal transmission techniques, RAG can significantly improve the generation quality.
@@
Instead of updating model parameters, prompt engineering utilizes LLM’s inference capabilities, designing specific input prompts to guide the generation of pre-trained LLMs. Importantly, prompt engineering avoids the need for backward passes and gradient updates of LLMs, and therefore the time costs are significantly lower than fine-tuning. Due to the low costs, it can quickly adapt to dynamic network environments, handling various network tasks with much lower response time. In addition, prompting engineering also aligns well with several crucial wireless network features. Firstly, many network devices are computational and energy resource-constrained, and prompt engineering has low requirements for computational resources, which may mitigate the energy consumption and computational burden on network devices. Secondly, prompt engineering allows human language-based instructions, e.g., network-specific task descriptions, questions, and solutions. By integrating natural language, network intelligence will become more accessible, especially for operators with minimal professional AI knowledge. In addition, prompt engineering has high design flexibility. With proper textual templates, prompts can be efficiently designed using human language and applied to various network tasks, which also have minimum requirements on mathematical equations. Moreover, prompting can be easily combined with other techniques, e.g., fine-tuning LLMs on network-specific datasets, and then prompting the model to further improve the generated content.
On the other hand, prompt engineering also requires considerable expertise in crafting the prompts, which will directly affect the quality of generated content. Although prompt engineering can significantly improve LLM’s performance, it relies on the model’s inherent capabilities, e.g., achieved by pre-training or fine-tuning. Therefore, given the opportunities and challenges, it is crucial to explore the integration of prompting techniques and wireless networks, showing a promising way to utilize generative AI.
@@
TABLE II: Summary of different prompting techniques for wireless networks.  (Updated)
@@
Prompting techniques Key principles Advantages Potential issues Wireless network applications
@@
In-context learning In-context learning empowers LLMs to learn from task-specific instruction and demonstrations. These prompts are designed with task-specific examples that are comprehensible to LLMs. LLMs can be generalized to perform various learning tasks by utilizing existing knowledge from pre-training data and acquiring new task-solving strategies from contextual demonstration prompts. Demonstrations greatly impact in-context learning, which should be carefully designed, e.g., selection, format, and order of examples. In-context learning is a promising approach to applying LLMs to wireless networks, using previously accumulated solutions as input prompts to address new network issues.
@@
Chain-of-thought 1) Chain-of-thought is an advanced strategy to improve LLM’s capabilities in complex reasoning. 2) It can guide LLMs to better understand the logical steps from questions to answers. LLMs can solve complex problems more transparently and logically through chain-of-thought demonstration prompts, e.g., “Let’s think step by step. Chain-of-thought prompting faces issues like incorrect reasoning and instability. Their inherent instability may provide incorrect and misleading intermediate reasoning steps. Chain-of-thought can make LLMs better adapt to complicated wireless environments that require step-by-step reasoning and planning, e.g., project code generation with multi-step scheduling.
@@
Prompt-based planning Prompt-based planning aims to address more complicated scenarios such as multi-hop question answering. It breaks complex tasks down into manageable sub-tasks. LLMs can make complex tasks more manageable by using their planning capabilities, lowering the overall difficulties. LLM-based planning requires dedicated analyses to decompose a complicated task into multiple sub-tasks. Zero-shot and automated task decomposition is still the main difficulty. Wireless networks are complicated systems and LLM-enabled automated task decomposition can be extremely useful for long-term network management and project development.
@@
Self-refinement Self-refinement allows them to correct their outputs through iterative feedback and refining demonstration prompts. Through iterative feedback generation and prediction refinement, LLMs can self-enhance their outputs with strong effectiveness and scalable capabilities. All historical outputs are iteratively appended as the inputs of LLMs, and it may lead to increasingly lengthy demonstration prompts and unavoidable resource consumption. Self-refinement will significantly save human efforts for network tasks, since LLMs can constantly improve the previous network policy automatically.